A pilot points his plane due north and flies at a speed with respect to air given by vpa=475 km/hr. A wind is blowing at vag=58 km/hr with respect to the ground at an angle of 50º south of east. What is the speed of the plane with respect to the ground?A pilot points his plane due north and flies at a speed with respect to air given by Vpa=475 km/hr. A wind is blowing at Vag=
Answer #1

Vpg = Vpa + Vag

in x - direction

Vpg,x = 0 + 58 * cos 50 = 37.28 km/h

in y-direction

Vpg,y = 475 - 58 * sin 50 = 430.57 km/h

so,

Vpg = sqrt ( 37.282 + 430.572)

Vpg = 432.2 Km/h